OpenStudy (anonymous):

Write the equation of a circle with center at (4, 1) and a radius of 6. A. (x - 4)2 + (y – 1)2 = 36 B. (x + 4)2 + (y + 1)2 = 36 C. (x - 4)2 + (y – 1)2 = 6 D. (x + 4)2 + (y + 1)2 = 6

OpenStudy (anonymous):

@robz8 how did you get that last one? i totally forgot to ask.

OpenStudy (anonymous):

okay well your center is (4, 1) where x = 4 and y = 1 so when you look at these choices, your parenthesis need to equal 0 so you're looking for (x-4) and (y-1) because (4-4) and (1-1) both equal 0 and all that = r^2 so your radius is 6 so your equation needs to equal 36 looks like A to me

OpenStudy (dls):

\[\Huge (x-h)^2+(y-x)^2=a^2\] is the equation of the circle

OpenStudy (anonymous):

Write the equation of a circle with center at (0, 0) and a diameter of 20. okay so for this one, is it just 20 since the center is 0, 0 ??

OpenStudy (dls):

\[\Huge x^2+y^2=a^2\]

OpenStudy (dls):

since -h and -k are 0

OpenStudy (anonymous):

^

OpenStudy (anonymous):

@chicagoblkhaks since your DIAMETER is 20, your radius would be 10 and remember, it equals the radius SQUARED to it would just be x^2 + y^2 = 100 since (x-0)^2 + (x-0)^2 = 10^2

OpenStudy (anonymous):

@DLS a little typo in your first equation btw :P

OpenStudy (dls):

oh lol sorry

OpenStudy (anonymous):

ohh. i knew that. hah. oops. 20 isnt even an option for me, i just typed it i guess thinking about the diameter. ugh. embarrassing